A 95N force is applied to a 412kg mass. How much does the mass accelerate ?
1
Expert's answer
2018-05-25T12:56:08-0400
We can find how much does the mass accelerate from the Newton’s Second Law of Motion:
F=ma,
here, F is the force applied to the mass, m is the mass and a is the acceleration of the mass.
Then, from this formula we can find the acceleration of the mass:
a=F/m=(95 N)/(412 kg)=0.23 m/s^2 .